# Shift current in 2D Janus Transition-Metal Dichalcogenides: the role of excitons

### v1

@@ Abstract
We study the shift current in two two-dimensional (2D) Janus transition metal dichalcogenides: molybdenum diselenide (MoSSe) and tungsten diselenide (WSSe). The shift current is evaluated using a real-time approach, in which the coupling with an external field is described in terms of a dynamical Berry phase. This approach incorporates electron-hole interactions and quasiparticle band structure renormalization through an effective Hamiltonian derived from many-body perturbation theory. We find that the shift current is strongly enhanced in correspondence of C excitons. An analysis in terms of the electron-hole pairs reveals that electron and hole are localized on different atoms, and thus following an optical excitation, the center of the electron charge is shifted thus giving rise to a significant photocurrent. These results highlight the role played by excitons in the shift-current response of Janus TMDs and demonstrate that these materials are promising building blocks for future photovoltaic devices.
@@ I Introduction
The shift current(SC) is a bulk photovoltaic effect that originates from the nonlinear interaction between light and a noncentrosymmetric crystal when illuminated by a continuous wave. As a second-order nonlinear phenomenon, SC has received significant attention in recent years.Cook et al. (2017) It involves the creation of a direct current (DC) within non-centrosymmetric materials when these are illuminated by light, demonstrating their intrinsic potential for application in solar cells. Although SC is a promising mechanism for generating photocurrent without the need for p–n junctions, the short-circuit current—which is closely linked to the power efficiency of a solar cell—is typically low. Therefore, it is crucial to understand the underlying physics of this phenomenon and to identify materials with high SC conductivity.
In a recent work, Cook et al.Cook et al. (2017) put forward the designing principles for optimizing the SC. They found that low dimensional systems can display a very strong SC due to a strong density of states due Van Hove singularities. Besides the density of states, in low-dimensional materials, due to the confinement of electrons and the poor screening of electron-electron interaction, the formation of bound states, the excitons, occurs upon the excitation of the material, which further enhances the SC.Chan et al. (2021); Hipolito et al. (2016); Esteve-Paredes et al. (2025); Morimoto and Nagaosa (2016); Fei et al. (2020) The need for a computational approach including all relevant many-body effects and specifically excitonic effects to predict the SC in materials is testified by the large disagreement between theory and experiments. On the one hand, so far most calculations are carried out at the independent particle level. On the other hand, in experiments there are extrinsic factors, difficult to control and quantify, that may modify the SC. For example in WS2 nanotubes, a strong SC response was reported deviating significantly from the theoretical prediction.Zhang et al. (2019) It has been argued that this discrepancy is due either to tube-tube interactions or the different cross section during the evaluation of SC.Krishna et al. (2023); Kim et al. (2022) Yet, the SC in these materials has been calculated neglecting excitonic effects that can, in principle, strongly enhance optical response, and thus the SC, in low-dimensional systems.
@@
In this work, we present a real-time approach for predicting the SC conductivity from first-principles in which the coupling between electrons and light is described in terms of a dynamical Berry phase. The Hamiltonian describing the system is derived from the many-body perturbation theory, which allows us to include correlation effects in both the electronic band structure (see, for example, Aryasetiawan et al.Aryasetiawan and Gunnarsson (1998)) and the response functions (see, for example, StrinatiStrinati (1988)). We have already applied this approach to predict other nonlinear optical effects, such as second harmonic generation, two-photon absorption, and sum-frequency generation.Grillo et al. (2024); Attaccalite et al. (2018); Pionteck et al. (2025) In those cases, we extracted nonlinear susceptibilities from the time-dependent polarization density. Here, we extend the approach to extract nonlinear conductivities from the polarization current.
@@
As an application, we calculate the SC conductivity of two Janus transition metal dichalcogenides (TMDs): MoSSe and WSSe, among the recently synthesized Janus 2D TMDs.Lu et al. (2017); Zhang et al. (2017); Lin et al. (2020) These materials have attracted a lot of attention from the scientific community due to their peculiar properties. Janus monolayers are a new type of 2D material that breaks out-of-plane symmetry and has a large built-in electric field. Janus 2D TMDs exhibit several interesting properties, such as a strong piezoelectric response, long exciton lifetime and enhanced interlayer coupling due to their intrinsic electric field.Torun et al. (2023) Their internal electric field renders Janus material ideal for photocatalysis Xia et al. (2018), and photovoltaic applications Aparicio-Huacarpuma et al. (2025). Their stacking-dependent dipole moment can be used to tune their intrinsic properties.Torun et al. (2023) Similarly to other nonlinear optical effects, geometrical symmetries play an important role in determining the SC. With respect to standard TMDs, Janus TMDs lack the $z\rightarrow-z$, symmetry and have an out-of-plane dipole allowing for a SC along the $z$-directionStrasser et al. (2022); Chen et al. (2023). Furthermore, swapping S with Se atoms switches the chirality of these structures and heterostructures can be designed to enhance the photocurrent.Strasser et al. (2022). Regarding the in-plane symmetries (xy-plane), though there are non zero elements of the second-order conductivity tensor, no net SC results for unpolarized light—such as sunlight. It has been shown, that a significant SC can be obtained by breaking the C3v symmetry either by strain, interaction with a substrateHung et al. (2023); Chen et al. (2023), or by forming nanotubes.Yang et al. (2025a)
In this work, we consider monolayer MoSSe and WSSe with light polarized along the $y$-direction and obtain the non zero component of the second-order conductivity tensor $\sigma^{(2)}_{y;yy}$ [^1] This component is equal, minus a sign, to the other non-zero in-plane components. At difference with previous works, Chen et al. (2023) we include the effect of electron-hole interactions. Section II presents the theoretical approach and computation details. Section III.1 shows the results for the linear optical response and the SC. Excitonic effects strongly redistribute the intensity and enhance the SC. From the analysis of the excitonic state in terms of the electronic structure, we explain the difference between the A/B and C excitons in the SC enhancement.
@@ II Theoretical methods
We present the definition of the SC response (Sec. II.1) and the real-time approach used to calculate it, as well as the procedures to extract nonlinear conductivity coefficients from the real-time current (Sec. II.2). Then, we detail all the computational parameters used in the simulations (Sec. II.3). As these are the first calculations of the SC with the Yambo codeSangalli et al. (2019), in Appendix A we detail the numerical implementation of the current operator. For validation, we compare with existing results on the GeS monolayerChan et al. (2021); Ibañez Azpiroz et al. (2018); Esteve-Paredes et al. (2025); Gomes et al. (2016); Cunningham et al. (2018) (see Supp. Mat.).
@@ II.1 The shift current
where the summation on $\boldsymbol{\omega}$ indicates the sum over all distinct $\omega_{1},\omega_{2}$ pairs. With $\omega_{1},\omega_{2}\neq 0$, the lowest order that produces DC current is the second order when $\omega=\omega_{1}+\omega_{2}=0$Sipe and Shkrebtii (2000) In case of a monochromatic field,
@@ II.2 Equation of motion
We simulate a system under the influence of an external, time-dependent electric field by solving an effective Schrödinger equation.
$i\hbar\frac{d}{dt}|v_{{\bf k},m}\rangle=\left(\hat{H}_{\bf k}^{0}+i{\bf\cal{E}}\cdot\partial_{\bf k}\right)|v_{{\bf k},m}(t)\rangle.$ (3)
where $|v_{{\bf k},m}(t)\rangle$ are the time-dependent valances states and the dipole operator has been replaced with ${\bf r}=i\partial_{\bf k}$ and the k-derivative has been performed using a covariant approach.Souza et al. (2004); Attaccalite and Grüning (2013) In the following we will omit the time-dependence of $|v_{{\bf k},m}(t)\rangle$ in order to simplify the notation. from the real-time dynamics of the valence states we calculate the current as:
@@
The derivative in ${\bf k}$ space is evaluated by means of finite-difference covariant formulation (see Souza et al. in Ref. Souza et al., 2004):
~$\tilde{\partial}_{\bf k}|v_{{\bf k}n}\rangle$ $\simeq$ $\frac{4}{3}\frac{|\tilde{v}_{{\bf k}+\Delta{\bf k},n}\rangle-|\tilde{v}_{{\bf k}-\Delta{\bf k},n}\rangle}{2\Delta{\bf k}}$ (5)
@@
where the tilde indicates that we used the covariant derivative in order to deal with the arbitrary phase factor associated with wave functions at different ${\bf k}$-points.Souza et al. (2004); Attaccalite and Grüning (2013) In Appendix A, we describe the numerical implementation of Eqs. (4) and (5) in the Yambo code Sangalli et al. (2019).
In Eq (3), the effective Hamiltonian is chosen depending on the level of accuracy we aim at, from independent-particle approximation (IPA) to the time-dependent adiabatic GW (TD-aGW).Attaccalite et al. (2011) The latter approximation is equivalent to the GW+BSE method in the linear response limit.Attaccalite et al. (2011); Strinati (1988)
To obtain the second order conductivity corresponding to the SC, we perturb the system with a sinusoidal electric field $E(t)=E_{0}\theta(t-t_{0})sin(\omega_{0}t)$, where $E_{0},\omega_{0}$ are the strength and frequency of the applied electric field and $\theta$ the Heaviside step function. Further, we add a term to the Hamiltonian to dephase the eigenmodes that are excited upon the field switch-on at $t_{0}$.Attaccalite and Grüning (2013) After transient effects from the field switch-on have decayed, the final current can be written as a Fourier series expansion: $J(t)=\sum_{j=-N}^{N}c_{j}e^{\mathrm{i}j\omega_{0}t}$. The coefficients $c_{j}$ can be obtained by a discrete Fourier transform by sampling $J(t)$ on a period $2\pi/\omega_{0}$. The components of the $n$-order conductivity tensor are directly related to the $c_{j}$ coefficients through powers of the applied electric field. This procedure is the analogue of that used to extract the nonlinear susceptibilities from the polarization density that has been detailed in Refs. Attaccalite and Grüning, 2013; Pionteck et al., 2025. In this case, we extract the conductivity corresponding to the SC from $c_{0}=2\sigma^{(2)}(0;\omega_{0},-\omega_{0})E_{0}^{2}+O(E_{0}^{4})$.
Figure 1: The electronic band structures of (a) MoSSe and (b) WSSe monolayers, at the Kohn-Sham level (blue solid lines) and including the quasiparticle corrections from $\mathrm{G_{0}W_{0}}$ calculations (red dashed lines).
@@ II.3 Computational details
The ground states of all the systems studied in this manuscript were calculated using Density Functional Theory (DFT) with the PBE functionals and the DOJO stringent pseudo-potential v4.1 to replace the core electrons.van Setten et al. (2018) We include also a van der Waals correction to the DFT functional.Grimme (2006) All DFT calculations were performed with the Quantum ESPRESSO code.Giannozzi et al. (2009) The linear and nonlinear optical response and quasiparticle corrections were calculated using the Yambo code Sangalli et al. (2019). For the quasiparticle correction, the G0W0 approximationAryasetiawan and Gunnarsson (1998) was applied, where the screened interaction was treated with the Godby-Needs plasmon-pole model (PPM).Stankovski et al. (2011)
System $\mathbf{k}$-points $N_{\mathrm{b}}$ $\epsilon_{\text{cut}}$(Ha) $\epsilon_{\text{bands}}$ $L_{z}$ (Å) $d_{\text{eff}}$ (Å)
MoSSe $24\times 24$ 19-26 5 400 36.30 6.15
WSSe $24\times 24$ 19-26 5 400 36.30 3.35
Table 1: The parameters used in the nonlinear response calculations for MoSSe and WSSe monolayers are: ${\bf k}$-point sampling; the range of bands for the response functions; the cut-off,$\epsilon_{\text{cut}}$, and the number of bands, $\epsilon_{\text{bands}}$ used to converge the dielectric function $\epsilon_{\bf G,\bf{G^{\prime}}}$, and expansion of the Green’s function in the GW calculations; the height of the supercell, $L_{z}$ and the effective layer thickness, $d_{\text{eff}}$.
We calculated the optical absorption at the independent particle approximation(IPA) on top of Kohn-Sham band structure, adding G0W0 corrections (IPA-GW) and including the local-field effects and electron-hole interaction by solving the Bethe-Salpeter equation in the Tamm-DancoffStrinati (1988) (GW+BSE). While the IPA and IPA-GW are solved within the linear response framework, the GW+BSE is formulated as an eigenvalue problem for the two-particle Hamiltonian:
@@
$+\left(f_{c\boldsymbol{k}}-f_{v\boldsymbol{k}}\right)\left(2\overline{V}_{\begin{subarray}{c}vc{\bf k}\\ v^{\prime}c^{\prime}{\bf k}^{\prime}\end{subarray}}-W_{\begin{subarray}{c}vc\boldsymbol{k}\\ v^{\prime}c^{\prime}\boldsymbol{k}^{\prime}\end{subarray}}\right)$ (6)
where $E_{n\boldsymbol{k}}$ are quasiparticle energies and $f_{n\boldsymbol{k}}$ the electronic occupations and $n$ run over the valence ($v$) and conduction bands ($c$). $\overline{V}$ is the Coulomb potential derived from the variation of the Hartree term and $W$ the screened electron-hole interaction derived from the screened exchange.Strinati (1988) The eigenvectors $A^{\lambda}_{cv\boldsymbol{k}}$ and eigenvalues $E_{\lambda}$ of the Hamiltonian in Eq. II.3 are respectively the excitonic wave-functions and energies. These quantities are used to build up the macroscopic dielectric function $\epsilon_{M}$ which writes:
@@
where $T_{\lambda}=\sum_{cv{\bf k}}d_{cv{\bf k}}A^{\lambda}_{cv{\bf k}}$ are the excitonic dipoles, $d_{cv{\bf k}}$ are the dipole matrix elements between the Kohn-Sham states with $d_{cv{\bf k}}=\langle v{\bf k}|\hat{r}|c{\bf k}\rangle$, and $i\eta$ with $\eta$ chosen to be 0.1 eV is a small broadening factor to simulate the experimental spectra. In order to consider isolated monolayers, a cutoff was applied to the Coulomb potential to avoid interaction between periodic replicas along the $z$-direction (Rozzi et al. Rozzi et al., 2006).
The nonlinear response is evaluated from the solution of the equation of motions as in Sec. II.2 at the corresponding levels of theory as for the optical absorptionAttaccalite et al. (2011): TD-IP, TD-IP@GW, and TD-aGW. The equations of motion have been solved using the Crank-Nicolson solverAttaccalite and Grüning (2013), with an integration time of 83 fs and a time-step of 0.01 fs and a dephasing time corresponding to a 0.1 eV spectral broadening. The nonlinear response was extracted from the real-time current using the YamboPy codePaleari et al. (2025). All other relevant computational details are presented in Table 1. The spin-orbit coupling was not included in the calculations, as it has been shown to have little effect on the nonlinear response of Janus TMDs Strasser et al. (2022).

### v2

@@ Abstract
We investigate the shift current in two-dimensional (2D) Janus transition-metal dichalcogenides (TMDs). The shift current is evaluated using a real-time approach, where the coupling with an external field is described in terms of a dynamical Berry phase. This methodology incorporates electron-hole interactions and quasiparticle band structure renormalization through an effective Hamiltonian derived from many-body perturbation theory. We find that the shift current is strongly enhanced in correspondence with C excitons. An analysis in terms of the electron-hole pairs reveals that electron and hole are localized on different atoms, and thus, following an optical excitation, the center of the electron charge is displaced, giving rise to a significant photocurrent. Janus TMDs, with their intrinsic out-of-plane asymmetry and tunable electronic properties, are particularly appealing for next-generation optoelectronic and energy-harvesting technologies. These results highlight the role of excitons in the shift-current response of Janus TMDs and demonstrate their potential as promising building blocks for future photovoltaic devices.
@@ I Introduction
The shift current (SC) is a direct current generated in non-centrosymmetric materials under continuous wave illumination, a second-order bulk photovoltaic effect arising from the nonlinear light-matter interaction. This phenomenon presents a promising mechanism for next-generation solar energy conversion, as it generates a photocurrent without the need for p-n junctions. However, the practical application of SC is hindered by typically low short-circuit currents, which are intrinsically linked to solar cell power conversion efficiency. Overcoming this limitation requires a deeper understanding of the underlying physics and the identification of materials with high SC conductivity.
In a recent work, Cook et al.[cook2017design] put forward designing principles for optimizing the SC. They found that low dimensional systems can display a very strong SC due to van Hove singularities leading to a strong density of states. This theoretical prediction is supported by experiments, such as the observation of a strong SC in WS2 nanotubes.[zhang2019enhanced] However, a significant discrepancy persists between these experimental measurements and standard theoretical predictions.[zhang2019enhanced] This disagreement has been attributed to extrinsic factors like tube-tube interactions or incorrect measurement of the tube cross-sections.[PhysRevB.108.165418, kim2022giant] Crucially, most theoretical models overlook a key intrinsic factor: strong electron-hole interactions.
In low-dimensional materials, reduced dielectric screening and quantum confinement lead to the formation of bound excitons, which are known to dramatically reshape the linear and nonlinear optical response.[chan2021giant, hipolito2016nonlinear, esteve2025excitons, morimoto, fei2020shift] Excitonic effects could, in principle, strongly modify—and potentially enhance—the SC, yet they are absent from most state-of-the-art calculations.[PhysRevB.108.165418] The lack of these essential many-body effects is likely a primary reason for the theory-experiment gap, underscoring the critical need for a computational approach that incorporates excitonic interactions to accurately predict the SC.
@@
In this work, we address this challenge by presenting a first-principles real-time approach to calculate the SC conductivity. Our method describes the light-matter interaction via a dynamical Berry phase and is built upon a Hamiltonian derived from many-body perturbation theory. This framework allows us to include quasiparticle corrections to the electronic band structure[aryasetiawan1998gw] and to incorporate both local-field effects and electron-hole interactions in the response functions.[strinati] We have previously successfully applied this methodology[attaccalite2013nonlinear, gruning2014second] to other nonlinear optical phenomena, including second harmonic generation and sum-frequency generation (see e.g. [grillo2024tunable, attaccalite2018two, pionteck2025sum]). Here, we extend it to compute nonlinear conductivities directly from the real-time polarization current.
@@
We apply this approach to investigate the SC in two Janus transition metal dichalcogenides (TMDs): MoSSe and WSSe.[lu2017janus, acsnano.7b03186, acsnano.9b10196] These two Janus TMDs have been recently synthesized in the form of monolayers and nanotubes.[lu2017janus, acsnano.7b03186, acsnano.9b10196] Janus monolayers are a novel class of 2D materials where the broken out-of-plane symmetry creates a large built-in electric field,[acs.nanolett.2c04724] leading to unique properties like a strong piezoelectric response, long exciton lifetimes.[acs.nanolett.2c04724] Their internal electric field renders Janus materials ideal for photocatalysis [PhysRevB.98.165424], and their stacking-dependent dipole moment can be used to tune these intrinsic properties.[acs.nanolett.2c04724] This $z$-symmetry breaking also makes them ideal candidates for the SC, as it allows for new tensor components, including an out-of-plane response.[strasser2022nonlinear, cpl_40_8_087201] Furthermore, swapping S with Se atoms switches the chirality of these structure and this property can be used to design heterostructures that enhance the photocurrent.[strasser2022nonlinear] In this work we focus specifically on the dominant in-plane tensor component, $\sigma^{(2)}_{yyy}$. Regarding the in-plane symmetries (xy-plane), though there are non zero elements of the second-order conductivity tensor, no net SC results for unpolarized light—such as sunlight. It has been shown, that a significant SC can be obtained by breaking the C3v symmetry either by strain, interaction with a substrate[hung2023nonlinear, cpl_40_8_087201], or by forming nanotubes.[yang2025janus]
The paper is organized as follows: in Section II we present our theoretical approach and give all computation details of our calculations; in Sec. III.1 results for MoSSe and WSSe are presented with an analysis of the excitonic state responsible for the SC enhancement; finally in Sec. IV we draw the conclusion and discuss further possible applications.
Our results reveal that excitonic effects cause a profound redistribution of the SC intensity: they strongly suppress the response at the A/B exciton resonance while dramatically enhancing it at the C-exciton resonance. By analyzing the real-space charge distribution of these excitons, we explain this stark contrast as due to the different spatial separation between electron and hole, leading to a substantial shift in the charge center upon photoexcitation—the fundamental driver of the SC.
With this work we not only resolve key aspects of the SC mechanism in 2D materials but also establish a robust computational framework for its accurate prediction, guiding the search for high-performance bulk photovoltaic materials.
@@ II Theoretical methods
We present the definition of the SC response (Sec. II.1) and the real-time approach used to calculate it, as well as the procedures to extract nonlinear conductivity coefficients from the real-time current (Sec. II.2). Then, we detail all the computational parameters used in the simulations (Sec. II.3). As these are the first calculations of the SC with the Yambo code[yambo], in Appendix A we detail the numerical implementation of the current operator. For validation, we compare with existing results on the GeS monolayer[chan2021giant, PhysRevB.97.245143, esteve2025excitons, PhysRevB.94.155428, PhysRevMaterials.2.034603] (see Supplemental Material).
@@ II.1 The shift current
where the summation on $\boldsymbol{\omega}$ indicates the sum over all distinct $\omega_{1},\omega_{2}$ pairs. With $\omega_{1},\omega_{2}\neq 0$, the lowest order that produces DC current is the second order when $\omega=\omega_{1}+\omega_{2}=0$[PhysRevB.61.5337] In case of a monochromatic field,
@@ II.2 Equation of motion
We simulate a system under the influence of an external, time-dependent electric field by solving an effective Schrödinger equation:
$i\hbar\frac{d}{dt}|v_{{\bf k},m}\rangle=\left(\hat{H}_{\bf k}+i{\bf\cal{E}}\cdot\partial_{\bf k}\right)|v_{{\bf k},m}(t)\rangle,$ (3)
where $\hat{H}_{\bf k}$ is the effective Hamiltonian, $|v_{{\bf k},m}(t)\rangle$ are the time-dependent valence states and the dipole operator has been replaced with ${\bf r}=i\partial_{\bf k}$ and the k-derivative has been performed using a covariant approach.[souza_prb, attaccalite2013nonlinear] In the following, we will omit the time-dependence of $|v_{{\bf k},m}(t)\rangle$ in order to simplify the notation. From the real-time dynamics of the valence states we calculate the current as:
@@
The derivative in $\bf k$} space is evaluated by means of finite-difference covariant formulation (see Souza et al. in Ref. souza_prb):
$\tilde{\partial}_{\bf k}|v_{{\bf k}n}\rangle$ $\simeq$ $\frac{4}{3}\frac{|\tilde{v}_{{\bf k}+\Delta{\bf k},n}\rangle-|\tilde{v}_{{\bf k}-\Delta{\bf k},n}\rangle}{2\Delta{\bf k}}$ (5)
@@
where the tilde indicates that we used the covariant derivative in order to deal with the arbitrary phase factor associated with wave functions at different $\bf k$}-points.[souza_prb, attaccalite2013nonlinear] In Appendix A, we describe the numerical implementation of Eqs. (4) and (5) in the Yambo code [yambo].
In Eq (3), the effective Hamiltonian is chosen depending on the level of accuracy we aim at, from independent-particle approximation (IPA) to the time-dependent adiabatic GW (TD-aGW).[attaccalite2011] The latter approximation is equivalent to the GW+BSE method in the linear response limit.[attaccalite2011, strinati]
To obtain the second order conductivity corresponding to the SC, we perturb the system with a sinusoidal electric field $E(t)=E_{0}\theta(t-t_{0})sin(\omega_{0}t)$, where $E_{0},\omega_{0}$ are the strength and frequency of the applied electric field, respectively, and $\theta$ is the Heaviside step function. Further, we add a term to the Hamiltonian to dephase the eigenmodes that are excited when the field switch-on at $t_{0}$.[attaccalite2013nonlinear] After transient effects have decayed, the final current can be written as a Fourier series expansion: $J(t)=\sum_{j=-N}^{N}c_{j}e^{\mathrm{i}j\omega_{0}t}$. The coefficients $c_{j}$ can be obtained by a discrete Fourier transform by sampling $J(t)$ on a period $2\pi/\omega_{0}$. The components of the $n$-order conductivity tensor are directly related to the $c_{j}$ coefficients through powers of the applied electric field. This procedure is the analogue of that used to extract the nonlinear susceptibilities from the polarization density that has been detailed in Refs. attaccalite2013nonlinear, pionteck2025sum. In this case, we extract the conductivity corresponding to the SC from $c_{0}=2\sigma^{(2)}(0;\omega_{0},-\omega_{0})E_{0}^{2}+O(E_{0}^{4})$.
Figure 1: The electronic band structures of (a) MoSSe and (b) WSSe monolayers, at the Kohn-Sham level (blue solid lines), including the quasiparticle corrections from $\mathrm{G_{0}W_{0}}$ calculations (red dots) and interpolated bands (red lines).. See Supplemental Material for the coordinates of the high-symmetry points used.
@@ II.3 Computational details
The ground states of all the systems studied in this manuscript were calculated using Density Functional Theory (DFT) with the PBE functionals and the DOJO stringent pseudo-potential v4.1 to replace the core electrons.[PseudoDojo] We include also a van der Waals correction to the DFT functional.[grimme2006semiempirical] All DFT calculations were performed with the Quantum ESPRESSO code.[pwscf] The linear and nonlinear optical response and quasiparticle corrections were calculated using the Yambo code [yambo]. For the quasiparticle correction, the G0W0 approximation[aryasetiawan1998gw] was applied, where the screened interaction was treated with the Godby-Needs plasmon-pole model (PPM).[stankovski2011g]
System $\mathbf{k}$-points $N_{\mathrm{b}}$ $\epsilon_{\text{cut}}$(Ha) $\epsilon_{\text{bands}}$ $L_{z}$ (a.u.) $d_{\text{eff}}$ (Å)
MoSSe $24\times 24$ 19-26 5 400 36.30 6.69
WSSe $24\times 24$ 19-26 5 400 36.30 6.49
[…]
We calculated the optical absorption at the independent particle approximation (IPA) on top of Kohn-Sham band structure, adding G0W0 corrections (IPA-GW) and including the local-field effects and electron-hole interaction by solving the Bethe-Salpeter equation in the Tamm-Dancoff[strinati] (GW+BSE). While the IPA and IPA-GW are solved within the linear response framework, the GW+BSE is formulated as an eigenvalue problem for the two-particle Hamiltonian:
@@
$+\left(f_{c\boldsymbol{k}}-f_{v\boldsymbol{k}}\right)\left(2\overline{V}_{\begin{subarray}{c}vc{\bf k}\\ v^{\prime}c^{\prime}{\bf k}^{\prime}\end{subarray}}-W_{\begin{subarray}{c}vc\boldsymbol{k}\\ v^{\prime}c^{\prime}\boldsymbol{k}^{\prime}\end{subarray}}\right),$ (6)
where $E_{n\boldsymbol{k}}$ are quasiparticle energies and $f_{n\boldsymbol{k}}$ the electronic occupations and $n$ run over the valence ($v$) and conduction bands ($c$). $\overline{V}$ is the Coulomb potential derived from the variation of the Hartree term and $W$ the screened electron-hole interaction derived from the screened exchange.[strinati] The eigenvectors $A^{\lambda}_{cv\boldsymbol{k}}$ and eigenvalues $E_{\lambda}$ of the Hamiltonian from Eq. II.3 are respectively the excitonic wave-functions and energies. These quantities are used to build up the macroscopic dielectric function $\epsilon_{M}$ which writes:
@@
where $T_{\lambda}=\sum_{cv{\bf k}}d_{cv{\bf k}}A^{\lambda}_{cv{\bf k}}$ are the excitonic dipoles, $d_{cv{\bf k}}$ are the dipole matrix elements between the Kohn-Sham states with $d_{cv{\bf k}}=\langle v{\bf k}|\hat{r}|c{\bf k}\rangle$, and $i\eta$ with $\eta$ chosen to be 0.1 eV is a small broadening factor to simulate the experimental spectra. In order to consider isolated monolayers, a cutoff was applied to the Coulomb potential to avoid interaction between periodic replicas along the $z$-direction (Rozzi et al. rozzi2006exact). Parameters used in the $G_{0}W_{0}$ and BSE (Tab. 1) are sufficient to have band gaps converged within 0.05 eV and exciton energies within 0.001 eV due to the error cancellation between band gap and excitonic binding energy (see also Sec. III in Supplemental Material). The $G_{0}W_{0}$ corrections used in the BSE and the real-time propagation are calculated directly on a regular grid using such convergence criteria. Exclusively for plotting, the $G_{0}W_{0}$ band structure is obtained by interpolation using smooth Fourier transform.[SKW]
The nonlinear response is evaluated from the solution of the equation of motion as in Sec. II.2 at the corresponding levels of theory as for the optical absorption[attaccalite2011]: TD-IP, TD-IP@GW, and TD-aGW. The equations of motion have been solved using the Crank-Nicolson solver[attaccalite2013nonlinear], with an integration time of 83 fs, a time-step of 0.01 fs and a dephasing time corresponding to a 0.1 eV spectral broadening. The nonlinear response was extracted from the real-time current using the YamboPy code[YamboPy]. All other relevant computational details are presented in Table 1. The spin-orbit coupling was not included in the calculations, as it has been shown to have little effect on the nonlinear response of Janus TMDs [strasser2022nonlinear].